# ProFit: Leveraging High-Value Signals in SFT via Probability-Guided Token Selection

### v1

@@ 2.1 Data-Efficient Instruction Tuning
Recent SFT adheres to the less is more principle (Zhou et al., 2023a; Zhang et al., 2025; Li et al., 2025b), evolving from complexity-based selection (Cao et al., 2023; Li et al., 2024a) to 2025’s importance-aware metrics like MIWV (Jiang et al., 2025) and ICL-based filtering (Wang et al., 2025a; Jiang et al., 2025). However, these coarse-grained sample-level methods treat pairs as atomic units, overlooking intra-sample low-information segments or stylistic noise (Pang et al., 2025; Qin et al., 2025), thus limiting the model’s focus on dense logical signals.
@@ 2.2 Token-Level Training Objectives
To address granularity limitations, research has pivoted to token-level optimization. While classical methods like Focal Loss (Lin et al., 2017) and Unlikelihood Training (Welleck et al., 2019) targeted hard tokens, recent LLM approaches like Rho-1 (Lin et al., 2024) and TIS-DPO (Liu et al., 2024) rely on costly external reference models. More efficient intrinsic methods like DFT (Wu et al., 2025b) employ probability-driven soft reweighting, and CFT (Ruan et al., 2025) further validates the need for supervising critical regions. While sharing the probability-driven motivation with DFT, ProFit diverges by adopting a strict hard masking strategy to efficiently filter out non-core expressions in a single step.
@@ 3.1 Preliminaries
LoRA is a parameter-efficient fine-tuning technique for LLMs (Hu et al., 2022; Wu et al., 2024). Its core premise is that the change in weights during model adaptation, denoted as $\Delta W$, possesses a low intrinsic rank. Let $W_{0}\in\mathbb{R}^{d\times k}$ denote the frozen pre-trained weight matrix. Instead of updating the full parameters, LoRA approximates the weight update $\Delta W$ by decomposing it into two low-rank matrices, $A$ and $B$:
@@
where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$ are low-rank matrices with rank $r\ll\min(d,k)$, and $\alpha$ is a scaling hyperparameter ensuring the stability of the adaptation.
@@ 4.1 Semantic Analysis
To accurately extract high-quality training signals, we performed a joint analysis of semantic importance and prediction probability for the tokens in multiple reference answers. Specifically, we utilize Gemini-3-pro as a semantic evaluator to annotate the tokens in the reference answers, classifying them into trivial tokens, which represent interchangeable stylistic variations, and core tokens, which encapsulate the essential reasoning logic. Subsequently, we used Qwen3-4B-Base (Yang et al., 2025a) to perform forward propagation and calculate the predicted probability for each token. This choice provides a computation-efficient yet capable proxy for language modeling, striking a balance between estimation quality and inference cost.
@@ 5.1 Experimental Setup
For the training data, we curated a subset of 2,000 samples from the BAAI-InfinityInstruct Dataset (Zhou et al., 2023a; Li et al., 2025a; Muennighoff et al., 2025), prioritizing high reward scores as done in Shadow-FT (Wu et al., 2025a). To comprehensively verify the effectiveness and generalization of our method, we conducted evaluations across a diverse set of LLMs, including the Qwen3 series (Yang et al., 2025a), Llama 3 series (AI@Meta, 2024), and OLMo 2 series (OLMo et al., 2024). We employed LLaMA-Factory (Zheng et al., 2024) as our primary training framework. All experiments are conducted on 8 H20 GPUs. For detailed hyperparameter settings, please refer to the Appendix B.
@@
  • Entropy-based tuning (Wang et al., 2025b): A selective strategy that updates parameters only on tokens with entropy.
@@ 6.4 Superior RL Initialization
We conduct experiments based on the Qwen3-0.6B-Base model. Specifically, we employ GRPO (Group Relative Policy Optimization, (Shao et al., 2024)) to optimize the models initialized with three different strategies: Base, DFT, and our proposed ProFit. We employ the DeepScaleR dataset as training prompts following Liu et al. (2025). All experiments are conducted on 32 H20 GPUs. For evaluation, in addition to the previously introduced MATH-500, we employ Minerva (Hendrycks et al., 2021) and OlympiadBench (He et al., 2024), which serve as comprehensive benchmarks covering tasks from general mathematical reasoning to diverse competition-level challenges.

### v2

@@ 2.1 Data-Efficient Instruction Tuning
Recent SFT adheres to the less is more principle (Zhou et al., 2023a; Zhang et al., 2025; Li et al., 2025c), evolving from complexity-based selection (Cao et al., 2023; Li et al., 2024a) to 2025’s importance-aware metrics like MIWV (Jiang et al., 2025) and ICL-based filtering (Wang et al., 2025a; Jiang et al., 2025). However, these coarse-grained sample-level methods treat pairs as atomic units, overlooking intra-sample low-information segments or stylistic noise (Pang et al., 2025; Qin et al., 2025), thus limiting the model’s focus on dense logical signals.
@@ 2.2 Token-Level Training Objectives
To address granularity limitations, research has pivoted to token-level optimization. While classical methods like Focal Loss (Lin et al., 2017) and Unlikelihood Training (Welleck et al., 2019) targeted hard tokens, recent LLM approaches like Rho-1 (Lin et al., 2024) and TIS-DPO (Liu et al., 2024) rely on costly external reference models. More efficient intrinsic methods like DFT (Wu et al., 2025b) employ probability-driven soft reweighting, and CFT (Ruan et al., 2025) further validates the need for supervising critical regions. Recently, Li et al. (2025a) theoretically validated via a model-capability continuum that suppressing low-probability tokens benefits strong-prior domains like math, supporting our masking strategy. While sharing the probability-driven motivation with DFT, ProFit diverges by adopting a strict hard masking strategy to efficiently filter out non-core expressions in a single step.
@@ 3.1 Preliminaries
LoRA is a parameter-efficient fine-tuning technique based on the premise that weight updates $\Delta W$ possess a low intrinsic rank (Hu et al., 2022; Wu et al., 2024). For a frozen weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA approximates the update via a low-rank decomposition $BA$, where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$ with $r\ll\min(d,k)$. The forward pass is modified as:
@@
where $\alpha$ is a scaling hyperparameter for adaptation stability.
@@ 4.1 Semantic Analysis
To accurately extract high-quality training signals, we performed a joint analysis of semantic importance and prediction probability for the tokens in multiple reference answers. Specifically, we utilize Gemini-3-pro (Google DeepMind, 2025) as a semantic evaluator to annotate the tokens in the reference answers, classifying them into trivial tokens, which represent interchangeable stylistic variations, and core tokens, which encapsulate the essential reasoning logic. Subsequently, we used Qwen3-4B-Base (Yang et al., 2025a) to perform forward propagation and calculate the predicted probability for each token. This choice provides a computation-efficient yet capable proxy for language modeling, striking a balance between estimation quality and inference cost.
@@ 5.1 Experimental Setup
For the training data, we curated a subset of 2,000 samples from the BAAI-InfinityInstruct Dataset (Zhou et al., 2023a; Li et al., 2025b; Muennighoff et al., 2025), prioritizing high reward scores as done in Shadow-FT (Wu et al., 2025a). To comprehensively verify the effectiveness and generalization of our method, we conducted evaluations across a diverse set of LLMs, including the Qwen3 series (Yang et al., 2025a), Llama 3 series (AI@Meta, 2024), and OLMo 2 series (OLMo et al., 2024). We employed LLaMA-Factory (Zheng et al., 2024) as our primary training framework. All experiments are conducted on 8 H20 GPUs. For detailed hyperparameter settings, please refer to the Appendix B.
@@
  • Entropy-based tuning (Wang et al., 2025b): A selective strategy that updates parameters only on tokens with high entropy.
@@ 6.4 Superior RL Initialization
We conduct experiments based on the Qwen3-0.6B-Base model. Specifically, we employ GRPO (Group Relative Policy Optimization (Shao et al., 2024)) to optimize the models initialized with three different strategies: Base, DFT, and our proposed ProFit. We employ the DeepScaleR dataset as training prompts following Liu et al. (2025). All experiments are conducted on 32 H20 GPUs. For evaluation, in addition to the previously introduced MATH-500, we employ Minerva (Hendrycks et al., 2021) and OlympiadBench (He et al., 2024), which serve as comprehensive benchmarks covering tasks from general mathematical reasoning to diverse competition-level challenges.